OpenStudy (anonymous):

P4 + 6Cl2 --> 4PCl3 I have 1 mol Cl2 and 2 mol P4 How do I predict the product? Which mol ratios do I use?

OpenStudy (anonymous):



OpenStudy (aaronq):

You would need to find the limiting reagent first. This can be done by dividing the moles of each reactant by their corresponding stoichiometric coefficient in the balanced equation: \(\sf \dfrac{moles~of~Cl_2}{6}=\dfrac{1}{6}mol\) \(\sf \dfrac{moles~of~P_4}{1}=\dfrac{2}{4}=\dfrac{1}{2}mol\) Now by comparison we see that \(Cl_2\) is the limiting reagent (less moles). We use the those moles and the same concept, make a ratio of the \(Cl_2\), \(PCl_3\) and their coefficients: \(\sf \dfrac{moles~of~Cl_2}{6}=\dfrac{moles~of~PCl_3}{4}\) \(\sf \dfrac{1~mol}{6}=\dfrac{moles~of~PCl_3}{4}\) \(\sf moles~of ~PCl_3=\dfrac{1*4}{6}mol=0.667 ~mol\)